OpenStudy (anonymous):

Brandon uses2/3 of a jar of peanut butter each week. He wants to know how many weeks it will take to use 4 jars. Use the model to find 4 ÷2/3 , the number of weeks it will take

OpenStudy (anonymous):

can one help me

OpenStudy (dumbcow):

rewrite 4 as 12/3 6*2/3 = 12/3 it takes 6 weeks

OpenStudy (anonymous):

ok thanks

OpenStudy (dumbcow):

Also note, when dividing by a fraction, multiply by the reciprocal --> 4* 3/2 = 6

OpenStudy (anonymous):

k